# A Comprehensive Survey of Dataset Distillation

### v3

@@ 6 Performance Comparison
TABLE I: Performance comparison of different dataset distillation methods on different datasets. Abbreviations of GM, TM, DM are for gradient matching, trajectory matching, and distribution matching, respectively. Whole denotes the test accuracy in terms of the whole target dataset. The factorized DD methods are marked by .
@@ 6.1 Standard Benchmark
For a comprehensive comparison, we also present the test accuracy w.r.t. a random select search, coreset approach, and the original target dataset. For the coreset approach, the Herding algorithm [62] is employed in this survey due to its superior performance. Notably, for most of parametric DD methods, ConvNet [24, 63], which consists of three convolutional blocks, is the default architecture to distill the synthetic data; and all the test accuracy of distilled data are obtained by training the same architecture of ConvNet for a fair comparison.
The empirical comparison results associated with MNIST, FashionMNIST, SVHN, CIFAR-10/100, and Tiny ImageNet are presented in Table I. In spite of severe scale issues and few implementations on ImageNet-1K, we present the DD results of ImageNet-1K in Table II for completion. As shown in these tables, many methods are not tested on some datasets due to the lack of meaning for easy datasets or scalability problems for large datasets. We preserve these blanks in tables for a more fair and comprehensive comparison. To make the comparison clear, we use the orange number for the best of two nonfactorized DD methods and blue number for the best factorized DD methods in each category in Table I. In addition, we note the factorized DD methods with . Based on the performance comparison in the tables, we observe the followings:
@@
  • Dataset distillation methods outperforms random pick and coreset selection by large margins.
@@ 6.2 Cross-Architecture Transferability
The standard benchmark in Table I only shows the performance of DD methods in terms of the specific ConvNet, while the distilled dataset is usually used to train other unseen network architectures. Therefore, measuring the DD performance on different architecture is also important for more comprehensive evaluation. However, there is not a standard set of architectures to evaluate cross-architecture transferability of DD algorithms, and the choice of architectures are highly different in the DD literature.
@@
  • The distilled data are not always possess the best performance across different architectures, which underlines the importance of evaluation on multiple architectures.
  • The factorized DD methods (IDC and KFS) possess better cross-architecture transferasbility i.e., suffer a smaller accuracy drop when encounter unseen architectures.
@@ 7.1 Graph Data
Regarding a graph dataset of $\mathcal{T}=\{\mathbf{A},\mathbf{X},\mathbf{Y}\}$ with $N$ nodes, where $\mathbf{A}$, $\mathbf{X}$, and $\mathbf{Y}$ are the adjacency matrix, the node feature matrix, and the node labels, respectively, a graph neural network $\mathcal{G}$ is trained on $\mathcal{T}$ for node classification. Jin et al. [66] adopted the gradient matching framework to distill the graph $\mathcal{T}$ to a small synthetic graph $\mathcal{S}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ with $N^{\prime}$ nodes and $N^{\prime}\ll N$.
However, the number of parameters in the adjacency matrix $\mathbf{A}^{\prime}$ grows quadratically as $N^{\prime}$ increases, which greatly impedes the scalability of graph distillation. Considering the implicit correlation between the graph structure and node features, the authors employ a MLP-based model to predict the graph structure based on node features:
@@
However, for the graph classification problem where adjacency matrix is discrete, the conventional DD algorithms are no longer applicable. To distill synthetic graphs with discrete adjacency matrix $\mathbf{A}^{\prime}\in\{0,1\}^{|\mathbf{A}^{\prime}|}$, Jin et al. [67] formulate $\mathbf{A}^{\prime}$ as a probabilistic graph with Bernoulli distribution
@@ 8.4 Other Applications
Using the characteristic of capturing the essence of datasets, [89, 90] utilized dataset distillation for visual design. Cazenavette et al. [89] generated the representative textures by randomly cropping the synthetic images during the distillation process. In addition to extracting texture, Chen et al. [90] imposed the synthetic images to model the outfit compatibility through dataset distillation.
@@ 9.1 Challenges
DD with complex labels. The majority of DD techniques lie in the single-label classification task where its label space is an integer set. As for other challenging tasks such as detection [97] or segmentation [98], their label spaces locate in much higher dimensional spaces. Unlike DD with single-label classification that the specific number of synthetic images (IPC) are pre-assigned for each class, enumerated assignment is impractical when the label space is highly-dimensional. Therefore, to represent the rich label space, high-dimensional labels of synthetic data require cautious adjusting, which significantly hinders the DD efficiency especially when jointly optimized with the input data.
@@ 9.2 Future Directions
Despite aforementioned challenges existed in DD, there are still many promising directions that are worth exploring to (1) enhance dataset distillation performance and (2) develop trustworthy dataset distillation.
@@
To enhance adversarial robustness, Tsilivis et al. [106] and Wu et al. [107] employed dataset distillation to extract the information of adversarial examples and generate robust datasets. Then standard network training on the distilled robust dataset is sufficient to achieve satisfactory robustness to adversarial perturbations, which substantially saves computing resources compared to expensive adversarial training. For backdoor attack, Liu et al. [108] considered injecting backdoor triggers into the small distilled dataset. Because the synthetic data has a small size, direct trigger insertion is less effective and also perceptible. Therefore, these coworkers turned to insert triggers into the target dataset during the dataset distillation procedure. In addition, they propose to iteratively optimize the triggers during the distillation process to preserve the triggers’ information for better backdoor attack performance. These defense and attack algorithms in terms of DD invoke the distillation of robust synthetic datasets, and models trained on them can better defend against various attacks.

### v4

@@ 6 Performance Comparison
TABLE I: Performance comparison of different dataset distillation methods on different datasets. Abbreviations of GM, TM, and DM are for gradient matching, trajectory matching, and distribution matching, respectively. Whole denotes the test accuracy in terms of the whole target dataset. The factorized DD methods are marked by .
@@ 6.1 Standard Benchmark
For a comprehensive comparison, we also present the test accuracy w.r.t. a random select search, coreset approach, and the original target dataset. For the coreset approach, the Herding algorithm [62] is employed in this survey due to its superior performance. Notably, for most parametric DD methods, ConvNet [24, 63], which consists of three convolutional blocks, is the default architecture to distill the synthetic data; and all the test accuracy of distilled data are obtained by training the same architecture of ConvNet for a fair comparison.
The empirical comparison results associated with MNIST, FashionMNIST, SVHN, CIFAR-10/100, and Tiny ImageNet are presented in Table I. In spite of severe scale issues and few implementations on ImageNet-1K, we present the DD results of ImageNet-1K in Table II for completion. As shown in these tables, many methods are not tested on some datasets due to the lack of meaning for easy datasets or scalability problems for large datasets. We preserve these blanks in tables for a more fair and comprehensive comparison. To make the comparison clear, we use orange numbers for the best of two nonfactorized DD methods and blue numbers for the best factorized DD methods in each category in Table I. In addition, we note the factorized DD methods with . Based on the performance comparison in the tables, we observe the following:
@@
  • Dataset distillation methods outperform random pick and coreset selection by large margins.
@@ 6.2 Cross-Architecture Transferability
The standard benchmark in Table I only shows the performance of DD methods in terms of the specific ConvNet, while the distilled dataset is usually used to train other unseen network architectures. Therefore, measuring the DD performance on different architectures is also important for a more comprehensive evaluation. However, there is not a standard set of architectures to evaluate the cross-architecture transferability of DD algorithms, and the choice of architectures is highly different in the DD literature.
@@
  • The distilled data do not always possess the best performance across different architectures, which underlines the importance of evaluation on multiple architectures.
  • The factorized DD methods (IDC and KFS) possess better cross-architecture transferability, i.e., suffer a smaller accuracy drop when encountering unseen architectures.
@@ 7.1 Graph Data
Regarding a graph dataset of $\mathcal{T}=\{\mathbf{A},\mathbf{X},\mathbf{Y}\}$ with $N$ nodes, where $\mathbf{A}$, $\mathbf{X}$, and $\mathbf{Y}$ are the adjacency matrix, the node feature matrix, and the node labels, respectively, a graph neural network $\mathcal{G}$ is trained on $\mathcal{T}$ for node classification. Jin et al. [66] adopted the gradient matching approach to distill the graph $\mathcal{T}$ to a small synthetic graph $\mathcal{S}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ with $N^{\prime}$ nodes and $N^{\prime}\ll N$.
However, the number of parameters in the adjacency matrix $\mathbf{A}^{\prime}$ grows quadratically as $N^{\prime}$ increases, which greatly impedes the scalability of graph distillation. Considering the implicit correlation between the graph structure and node features, the authors employ an MLP-based model to predict the graph structure based on node features:
@@
However, for the graph classification problem where the adjacency matrix is discrete, the conventional DD algorithms are no longer applicable. To distill synthetic graphs with discrete adjacency matrix $\mathbf{A}^{\prime}\in\{0,1\}^{|\mathbf{A}^{\prime}|}$, Jin et al. [67] formulate $\mathbf{A}^{\prime}$ as a probabilistic graph with Bernoulli distribution
@@ 8.4 Other Applications
Leveraging the characteristic of capturing the essence of datasets, Cazenavette et al. [89], Chen et al. [90] adopted dataset distillation for visual design. Specifically, Cazenavette et al. [89] generated the representative textures by randomly cropping the synthetic images during the distillation process. In addition to extracting texture, Chen et al. [90] imposed the synthetic images to model the outfit compatibility through dataset distillation.
@@ 9.1 Challenges
DD with complex labels. The majority of DD techniques lie in the single-label classification task where its label space is an integer set. As for other challenging tasks such as detection [97] or segmentation [98], their label spaces are located in much higher dimensional spaces. Unlike DD with single-label classification that the specific number of synthetic images (IPC) are pre-assigned for each class, enumerated assignment is impractical when the label space is highly dimensional. Therefore, to represent the rich label space, high-dimensional labels of synthetic data require cautious adjusting, which significantly hinders the DD efficiency especially when jointly optimized with the input data.
@@ 9.2 Future Directions
Despite aforementioned challenges existing in DD, there are still many promising directions that are worth exploring to (1) enhance dataset distillation performance and (2) develop trustworthy dataset distillation.
@@
To enhance adversarial robustness, Tsilivis et al. [106] and Wu et al. [107] employed dataset distillation to extract the information of adversarial examples and generate robust datasets. Then standard network training on the distilled robust dataset is sufficient to achieve satisfactory robustness to adversarial perturbations, which substantially saves computing resources compared to expensive adversarial training. For backdoor attack, Liu et al. [108] considered injecting backdoor triggers into the small distilled dataset. Because the synthetic data possess a small size, direct trigger insertion is less effective and also perceptible. Therefore, these coworkers turned to insert triggers into the target dataset during the dataset distillation procedure. In addition, they propose to iteratively optimize the triggers during the distillation process to preserve the triggers’ information for better backdoor attack performance. These defense and attack algorithms in terms of DD invoke the distillation of robust synthetic datasets, and models trained on them can better defend against various attacks.